OpenStudy (anonymous):

x + y = k x - y = k The solution of the system shown is _____. (0, k) (0, 0) (k, 0) im not shure.help

OpenStudy (anonymous):

i think b

OpenStudy (anonymous):

If adding y and subtracting y lead to the same result, then y must be 0. So then x has to be...?

OpenStudy (anonymous):

I THINK B

OpenStudy (anonymous):

If x and y are both 0 then x + y = 0

OpenStudy (anonymous):

am i correct?????????

OpenStudy (anonymous):

I believe it is Zero. Because the equations are linear Graph am I correct? So X is going to be X for both Equations and so I so is Y and in this case X=0 and Y=0 so X + Y = K is Really 0 + 0 = 0 and X - Y = K is really 0 - 0 = 0 So Yes You are correct :D